OpenStudy (anonymous):

1/tanx + tanx=sec^2x/tanx

OpenStudy (solomonzelman):

1/tanx + tanx=sec^2x/tanx SAY: tanx= tan^2x/tanx this way you put everything under a denominator of tanx which would mean that all denominators of tanx cancel.

OpenStudy (solomonzelman):

1/tanx + tanx = sec^2x/tanx 1/tanx + tan^2x/tanx = sec^2x/tanx multiply everything by tanx, 1 + tan^2x = sec^2x identity proved.

OpenStudy (solomonzelman):

This only works if you were allowed to manipulate both sides, but if you were only able to work the one side, work the LEFT SIDE. 1/tanx + tanx = sec^2x/tanx 1/tanx + tan^2x/tanx = sec^2x/tanx add the fractions on the left 1 + tan^ 2x /tanx = sec^2x/tanx 1+tan^2x=sec^2x, so the left side becomes sec^2x/tanx . Hence is the proof, result is, sec^2x/tanx =sec^2x/tanx

OpenStudy (anonymous):

I dont understand how 1+tan^2x/tanx could equal 1+tan^2x, wouldn't the tan^2x being divided by tanx make it equal 1+tanx?

OpenStudy (solomonzelman):

Lets start over, okay?

OpenStudy (solomonzelman):

Tan x = Tan ^2x / Tanx right?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

so, I am substituting Tan ^2x / Tanx instead of Tan x. \[\frac{1}{Tanx}+\color{red}{Tanx }=\frac{Sec^2x}{Tanx} \] \[\frac{1}{Tanx}+\color{red}{ \frac{Tan^2x}{Tanx}} =\frac{Sec^2x}{Tanx} \] Now, I am adding fractions on the left side, \[\frac{1+Tan^2x}{Tanx} =\frac{Sec^2x}{Tanx} \]

OpenStudy (solomonzelman):

Knowing that \[\color{red}{ Tan^2x+1=Sec^2x}\] your identity is proved with a last substitution (substitute Sec^2x for 1+Tan^2x )

OpenStudy (anonymous):

OH! Wow okay, so obvious. I don't know why it I couldn't get it through my brain. Thank you!

OpenStudy (solomonzelman):

Becuase I didn't post it in latex. This latex just gives your eyes the chance to see it :)

OpenStudy (solomonzelman):

YW!

OpenStudy (anonymous):

would you mind helping me with one more?

OpenStudy (solomonzelman):

Sure.

OpenStudy (anonymous):

It's probably simple too, but it is csc(-x)/sec(-x)=-cot I got to -sin/cosx=-cot but Im not sure how to make what I have into the cotangent when it is just tangent right now

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{\csc(-x)}{\sec(-x)} =-\cot(x) } \] this?

OpenStudy (anonymous):

yes